OpenStudy (anonymous):

How can 1/6x − 5 = 1/5x + 2 be set up as a system of equations?

OpenStudy (anonymous):

answer choices are: 6y − x = −30 5y − x = 10 5y + x = −30 5y + x = 10 6y − 6x = −30 5y − 5x = 10 5y + 5x = −30 5y + 5x = 10

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

\[\frac{ 1 }{ 6 }x-5=y,x-5*6=6 y,~or~6 y-x=-30\] \[\frac{ 1 }{ 5 }x+2=y,x+2*5=5 y,or~5 y-x=10\]

OpenStudy (anonymous):

thank you! That's what I thought just needed and extra opinion.

OpenStudy (anonymous):

you are welcome.